OpenStudy (anonymous):

Find the volume of a sphere that has a surface area of 16 sq. in.

OpenStudy (anonymous):

area of sphere is 4(pi)(r^2). you know surface area, and everything is constant, but r. so get positive r from this eqn. Then you know that the volume is V= 4(pi)(r^3)/3, sub that positive r to get volume.

OpenStudy (anonymous):

Im confused...

OpenStudy (anonymous):

I know the area of a sphere but I don't understand the rest of what you said. haha sorry

OpenStudy (anonymous):

this question has one sub question in it. it is to get radius r from the given surface area. Surface area= 4(pi)r^2=16. get r.

OpenStudy (anonymous):

oh! ok! thank you for!

OpenStudy (anonymous):

after you have solved for r, you can get volume right?

OpenStudy (anonymous):

yes I believe the volume is \[\frac{ 32 }{ 3 }\] times pi

OpenStudy (anonymous):

:)